# Experimentation on Endogenous Graphs

### v1

@@ 5 Endogenous Unipartite Graphs
An endogenous unipartite interference graph is the special case of Definition 2, where the analysis and randomization units are the same: $\mathcal{A}=\mathcal{R}$. This definition allows for directed interference, in the sense that $E_{ar}$ may not equal $E_{ra}$. We specify the outcome model for all $a$ as $y_{a}=\alpha_{a}+\beta_{a}x_{a}+\gamma_{a}T_{a}$, where $x_{a}=\sum_{r}T_{r}e_{ar}w_{ar}$ and we impose $w_{aa}=0$. Thus $\beta_{a}x_{a}$ is the indirect effect on $a^{\prime}s$ outcome, coming from the treatment of others, and $\gamma_{a}T_{a}$ is the direct effect, coming from $a$ itself being treated. The TTE is $\frac{1}{n_{a}}\sum_{a}[\mathcal{W}_{a}(\mathbf{1})\beta_{a}+\gamma_{a}]$. Let $\hat{\gamma}_{a}=T_{a}y_{a}/p-(1-T_{a})y_{a}/(1-p)$. The previous result on unbiasedness in the bipartite setting can be directly applied to construct an class of unbiased estimators in the unipartite setting. We employ the previously developed bipartite estimator to estimate the indirect effect, and use $\hat{\gamma}_{a}$ as an unbiased estimator of the direct effect. Summing the two gives an unbiased estimator of the TTE, $\hat{\mu}^{u,c}_{uni}$, defined as
@@
The restriction $w_{aa}=u_{aa}=0$ is innocuous, given the presence of a separate term $\gamma_{a}T_{a}$ in the outcome model for the direct effect. The summands in (6) follow the same dependence structure as those in (4). Consistency and asymptotic normality consequently follow directly from the same arguments as the bipartite case, under the same conditions on the growth of the maximum degrees, $d_{\mathcal{A}}$ and $d_{\mathcal{R}}$.
@@ 6 Conclusion
The findings of this paper establish a foundation for estimating and testing treatment effects on endogenous graphs, but significant extensions and open questions remain. Future research could characterise the variance-minimizing choice of weights $u_{ar}$ used in the construction of the instruments $z^{u}_{a}$. A consistent and scalable variance estimator could be used in conjunction with Theorem 2 to construct confidence intervals around the TTE estimator, complementing the testing results of Section 4. And while we generalize the linear outcome model from the conventional exogenous edge setting, the outcome model may be implausible in some applications. We believe the research direction of causal inference on endogenous graphs holds promise for advancing our understanding of treatment effects in complex, real-world networks.
@@ Acknowledgements
Appendix A Proofs
@@
$\text{Cov}\left(\sum_{r}s_{ar}(T_{r})T_{r},\sum_{r}v_{ar}(T_{r})T_{r}\right)=p(1-p)\sum_{r}s_{ar}(1)v_{ar}(1),$
@@
$\text{Cov}\left(\sum_{r}s_{ar}(T_{r})T_{r},\sum_{r}v_{ar}(T_{r})T_{r}\right)$ $=\text{Cov}\left(\sum_{r}s_{ar}(1)T_{r},\sum_{r}v_{ar}(1)T_{r}\right)$
$=\sum_{r}\textrm{Cov}\left(s_{ar}(1)T_{r},v_{ar}(1)T_{r}\right)+\sum_{r}\sum_{r^{\prime}\neq r}\underbrace{\textrm{Cov}\left(s_{ar}(1)T_{r},v_{ar^{\prime}}(1)(T_{r^{\prime}})\right)}_{=0,\text{ pairwise independence}}$
@@
In the setting of Definition 2, suppose $G$ is an anchor sub-graph. If (a) edges are $r$-driven; (b) the treatment assignments are independent Bernoulli random variables with probability $p$; (c) $\{(a,r)\mid u_{ar}\neq 0\}\subset G$; (d) $c_{ar}=\mathbb{I}((a,r)\in G)$; (e) $w_{ar}\neq 0$ for all $(a,r)$; and (f) |$|\{r\mid u_{ar}\neq 0\}|>0$ for all $a$, then
@@
$\textrm{Cov}\left(\widehat{\mathcal{W}_{a}(\mathbf{1})},y_{a}\left(\frac{z_{a}-\mathbb{E}{z_{a}}}{\textrm{Cov}(x_{a},z_{a})}\right)\right)\\ =\underbrace{\mathbb{E}\left[\textrm{Cov}\left(\widehat{\mathcal{W}_{a}(\mathbf{1})},y_{a}\left(\frac{z_{a}-\mathbb{E}{z_{a}}}{\textrm{Cov}(x_{a},z_{a})}\right)\mid T_{A}\right)\right]}_{X}+\underbrace{\textrm{Cov}\left(\mathbb{E}\left[\widehat{\mathcal{W}_{a}(\mathbf{1})}\mid T_{A}\right],\mathbb{E}\left[y_{a}\left(\frac{z_{a}-\mathbb{E}{z_{a}}}{\textrm{Cov}(x_{a},z_{a})}\right)\mid T_{A}\right]\right)}_{Y},$
@@
$\textrm{Cov}\left(\widehat{\mathcal{W}_{a}(\mathbf{1})},y_{a}\left(\frac{z_{a}-\mathbb{E}{z_{a}}}{\textrm{Cov}(x_{a},z_{a})}\right)\mid T_{A}\right)$ $=\left(\frac{z_{a}-\mathbb{E}{z_{a}}}{\textrm{Cov}(x_{a},z_{a})}\right)\textrm{Cov}\left(\widehat{\mathcal{W}_{a}(\mathbf{1})},y_{a}\mid T_{A}\right)$
[…]
is the product of a constant and a mean-zero random variable, so $X$, its expectation, is zero.
@@
$\mathbb{E}\left[\widehat{\mathcal{W}_{a}(\mathbf{1})}\mid T_{A}\right]=\underbrace{\sum_{r\in A}T_{r}w_{ar}(e_{ar}-c_{ar})/p}_{=0}+\sum_{r\in A}w_{ar}c_{ar}+\sum_{r\not\in A}w_{ar}e_{ar}(1),$
where the first term is zero because $T_{r}w_{ar}(e_{ar}-c_{ar})=T_{r}w_{ar}(e_{ar}(1)-c_{ar})$, and $e_{ar}(1)=c_{ar}=1$ for $r\in A$. This makes $\mathbb{E}\left[\widehat{\mathcal{W}_{a}(\mathbf{1})}\mid T_{A}\right]$ an constant, and thus $Y=0$. ∎
@@
$=\sum_{r\in V_{a}}\left(\frac{T_{r}w_{ar}(e_{ar}(1)-1)}{p}+w_{ar}\right)+\sum_{r\not\in V_{a}}\left(\frac{T_{r}w_{ar}e_{ar}(1)}{p}\right)$
@@
$=\sum_{r\in V_{a}}w_{ar}+\sum_{r\in R_{a}\setminus V_{a}}\left(\frac{T_{r}w_{ar}e_{ar}(1)}{p}\right)+\sum_{r\not\in R_{a}}\left(\frac{T_{r}w_{ar}e_{ar}(1)}{p}\right)$
@@
$\text{Var}(\hat{\mu})$ $=\frac{1}{n_{a}^{2}}\sum_{i\in\mathcal{A}}\sum_{j\in\mathcal{A}}\text{Cov}(\hat{\mu}_{i},\hat{\mu}_{j})$ (11)
$=\frac{1}{n_{a}^{2}}\sum_{i\in\mathcal{A}}\sum_{j\in\mathcal{I}(i)}\text{Cov}(\hat{\mu}_{i},\hat{\mu}_{j})\quad\text{($\hat{\mu}_{i}$ is independent of $\hat{\mu}_{j}$ if $R_{i}\cap R_{j}=\emptyset$)}$
@@
$|a_{i}|\leq|\hat{\mu}_{i}^{u,c}|+|\mu_{i}|=|\hat{\mu}_{i}^{u,c}|+|\mathbb{E}[\hat{\mu}_{i}^{u,c}]|\leq|\hat{\mu}_{i}^{u,c}|+\mathbb{E}[|\hat{\mu}_{i}^{u,c}|]\leq\frac{2MW_{h}|R_{a}|}{p^{2}(1-p)W_{l}}\leq\frac{2MW_{h}d_{a}}{p^{2}(1-p)W_{l}}.$ (12)
@@
$d_{W}\left(\frac{\hat{\mu}^{u,c}-\mu}{\sqrt{\text{Var}(\hat{\mu}^{u,c})}},Z\right)\leq\frac{d_{a}^{2}d_{r}^{2}}{\sigma^{3}n_{a}^{2}}\left[\frac{2MW_{h}d_{a}}{p^{2}(1-p)W_{l}}\right]^{3}+\sqrt{\frac{28}{\pi}}\frac{d_{a}^{3/2}d_{r}^{3/2}}{\sigma^{2}n_{a}^{3/2}}\left[\frac{2MW_{h}d_{a}}{p^{2}(1-p)W_{l}}\right]^{2}\\ =O\left(\frac{d_{a}^{5}d_{r}^{2}}{\sigma^{3}n_{a}^{2}}+\frac{d_{a}^{7/2}d_{r}^{3/2}}{\sigma^{2}n_{a}^{3/2}}\right)=O\left(\frac{d_{a}^{5}d_{r}^{2}}{n_{a}^{1/2}}+\frac{d_{a}^{7/2}d_{r}^{3/2}}{n_{a}^{1/2}}\right)=O\left(\frac{d_{a}^{5}d_{r}^{2}}{n_{a}^{1/2}}\right).$
@@
Lemma 4 (Lemma 3.6 in Ross, (2011)).
@@
[…]
@@
$=\frac{1}{n_{a}^{2}}\sum_{i\in\mathcal{A}}\sum_{j\in\mathcal{J}(i)}\text{Cov}(\tilde{\mu}_{i},\tilde{\mu}_{j}\mid y)\quad\text{(if $V_{i}\cap V_{j}=\emptyset$, $z_{i}^{u}$ is independent of $z_{j}^{u}$ and thus $\tilde{\mu}_{i}$ is independent of $\tilde{\mu}_{j}$)}$
@@
$=\mathbb{E}[\tilde{\mu}^{2}\mid y,|\tilde{\mu}|\geq c_{\alpha}]P(|\tilde{\mu}|\geq c_{\alpha}\mid y)+\mathbb{E}[\tilde{\mu}^{2}\mid y,|\tilde{\mu}|<c_{\alpha}]P(|\tilde{\mu}|<c_{\alpha}\mid y)$
@@
As for $\text{Var}(\tilde{\mu})$,
$\text{Var}(\tilde{\mu})$ $=\frac{1}{n_{a}^{2}}\sum_{i\in\mathcal{A}}\sum_{j\in\mathcal{A}}\text{Cov}(\tilde{\mu}_{i},\tilde{\mu}_{j})$ (16)
$=\frac{1}{n_{a}^{2}}\sum_{i\in\mathcal{A}}\sum_{j\in\mathcal{I}(i)}\text{Cov}(\tilde{\mu}_{i},\tilde{\mu}_{j})\quad\text{(if $R_{i}\cap R_{j}=\emptyset$, $\tilde{\mu}_{i}$ is independent of $\tilde{\mu}_{j}$)}$
@@
[…]

### v2

@@ 6 Endogenous Unipartite Graphs
An endogenous unipartite interference graph is the special case of Definition 2, where the analysis and randomization units are the same: $\mathcal{A}=\mathcal{R}$. This definition allows for directed interference, in the sense that $E_{ar}$ may not equal $E_{ra}$. We specify the outcome model for all $a$ as $y_{a}=\alpha_{a}+\beta_{a}x_{a}+\gamma_{a}T_{a}$, where $x_{a}=\sum_{r}T_{r}e_{ar}w_{ar}$ and we impose $w_{aa}=0$. Thus, $\beta_{a}x_{a}$ is the indirect effect on $a^{\prime}s$ outcome coming from the treatment of others, and $\gamma_{a}T_{a}$ is the direct effect coming from $a$ itself being treated. The TTE is $\frac{1}{n_{a}}\sum_{a}[\mathcal{W}_{a}(\mathbf{1})\beta_{a}+\gamma_{a}]$. Let $\hat{\gamma}_{a}=T_{a}y_{a}/p-(1-T_{a})y_{a}/(1-p)$. The previous result on unbiasedness in the bipartite setting can be directly applied to construct a class of unbiased estimators in the unipartite setting. We employ the previously developed bipartite estimator to estimate the indirect effect and use $\hat{\gamma}_{a}$ as an unbiased estimator of the direct effect. Summing the two gives an unbiased estimator of the TTE, $\hat{\mu}^{u,c}_{uni}$, defined as
@@
The restriction $w_{aa}=u_{aa}=0$ is innocuous given the presence of a separate term $\gamma_{a}T_{a}$ in the outcome model for the direct effect. The summands in (6) follow the same dependence structure as those in (4). Consistency and asymptotic normality consequently follow directly from the same arguments as the bipartite case, under the same conditions on the growth of the maximum degrees, $d_{\mathcal{A}}$ and $d_{\mathcal{R}}$.
@@ 7 Conclusion
We establish a foundation for estimating and testing treatment effects on endogenous graphs, but significant extensions and open questions remain. Future research could characterize the variance-minimizing choice of weights $u_{ar}$ used to construct instruments $z^{u}_{a}$. A consistent and scalable variance estimator could be used in conjunction with Theorem 2 to construct confidence intervals around the TTE estimator, complementing the testing results of Section 4. And while we generalize the linear outcome model from the conventional exogenous edge setting, the outcome model may be implausible in some applications. We believe the research direction of causal inference on endogenous graphs holds promise for advancing our understanding of treatment effects in complex, real-world networks.
@@ Acknowledgements
Appendix A Variance and Confidence Interval for ERL Estimators
Let
$x_{a}^{\text{pre}}(\mathbf{T})=\frac{\sum_{r\in\mathcal{R}}e_{ar}^{\text{pre}}T_{r}}{\sum_{r\in\mathcal{R}}e_{ar}^{\text{pre}}}$
and
$x_{a}^{\text{post}}(\mathbf{T})=\frac{\sum_{r\in\mathcal{R}}e_{ar}^{\text{pre}}T_{r}}{\sum_{r\in\mathcal{R}}e_{ar}^{\text{post}}},$
where $e_{ar}^{\text{pre}}$ and $e_{ar}^{\text{post}}$ are the pre-treatment and post-treatment edge between $a$ and $r$. The ERL estimators (Harshaw et al. 2021) are defined as
$\hat{\mu}^{\text{pre}}=\frac{1}{n_{a}}\sum_{a}\frac{y_{a}(x_{a}^{\text{pre}}(\mathbf{T})-p)}{p(1-p)/(\sum_{r\in\mathcal{R}}e_{ar}^{\text{pre}})}$
and
$\hat{\mu}^{\text{post}}=\frac{1}{n_{a}}\sum_{a}\frac{y_{a}(x_{a}^{\text{pre}}(\mathbf{T})-p)}{p(1-p)/(\sum_{r\in\mathcal{R}}e_{ar}^{\text{post}})}$
In stead of using the variance estimator proposed in Harshaw et al. (2021), we calculate $\text{Var}(\hat{\mu}^{\text{pre}}\mid\mathbf{y},\mathbf{e}^{\text{pre}})$ and $\text{Var}(\hat{\mu}^{\text{post}}\mid\mathbf{y},\mathbf{e}^{\text{post}})$ under the strong null, where $(\mathbf{y},\mathbf{e}^{\text{pre}},\mathbf{e}^{\text{post}})$ is independent of $\mathbf{T}$, and use them to construct confidence intervals based on asymptotic normality in our simulations in Section 5.
Appendix B Proofs
@@
$\text{Cov}\left(\sum_{r}s_{ar}(T_{r})T_{r},\sum_{r}v_{ar}(T_{r})T_{r}\right)$
$=p(1-p)$ $\sum_{r}s_{ar}(1)v_{ar}(1),$
@@
$\quad\text{Cov}\left(\sum_{r}s_{ar}(T_{r})T_{r},\sum_{r}v_{ar}(T_{r})T_{r}\right)$
$=\text{Cov}\left(\sum_{r}s_{ar}(1)T_{r},\sum_{r}v_{ar}(1)T_{r}\right)$
$=\sum_{r}\textrm{Cov}\left(s_{ar}(1)T_{r},v_{ar}(1)T_{r}\right)$
$\quad{}+\sum_{r}\sum_{r^{\prime}\neq r}\underbrace{\textrm{Cov}\left(s_{ar}(1)T_{r},v_{ar^{\prime}}(1)(T_{r^{\prime}})\right)}_{=0,\text{ pairwise independence}}$
@@
In the setting of Definition 2, suppose $G$ is an anchor sub-graph. If (a) edges are $r$-driven; (b) the treatment assignments are independent Bernoulli random variables with probability $p$; (c) $\{(a,r)\mid u_{ar}\neq 0\}\subset G$; (d) $c_{ar}=\mathbb{I}((a,r)\in G)$; (e) $w_{ar}\neq 0$ for all $(a,r)$; and (f) $|\{r\mid u_{ar}\neq 0\}|>0$ for all $a$, then
@@
$\textrm{Cov}\left(\widehat{\mathcal{W}_{a}(\mathbf{1})},y_{a}\left(\frac{z_{a}-\mathbb{E}{z_{a}}}{\textrm{Cov}(x_{a},z_{a})}\right)\right)\\ =\underbrace{\mathbb{E}\left[\textrm{Cov}\left(\widehat{\mathcal{W}_{a}(\mathbf{1})},y_{a}\left(\frac{z_{a}-\mathbb{E}{z_{a}}}{\textrm{Cov}(x_{a},z_{a})}\right)\mid T_{A}\right)\right]}_{X}\\ +\underbrace{\textrm{Cov}\left(\mathbb{E}\left[\widehat{\mathcal{W}_{a}(\mathbf{1})}\mid T_{A}\right],\mathbb{E}\left[y_{a}\left(\frac{z_{a}-\mathbb{E}{z_{a}}}{\textrm{Cov}(x_{a},z_{a})}\right)\mid T_{A}\right]\right)}_{Y},$
@@
$\quad\textrm{Cov}\left(\widehat{\mathcal{W}_{a}(\mathbf{1})},y_{a}\left(\frac{z_{a}-\mathbb{E}{z_{a}}}{\textrm{Cov}(x_{a},z_{a})}\right)\mid T_{A}\right)$
$=\left(\frac{z_{a}-\mathbb{E}{z_{a}}}{\textrm{Cov}(x_{a},z_{a})}\right)\textrm{Cov}\left(\widehat{\mathcal{W}_{a}(\mathbf{1})},y_{a}\mid T_{A}\right)$
$=\left(\frac{z_{a}-\mathbb{E}{z_{a}}}{\textrm{Cov}(x_{a},z_{a})}\right)\underbrace{\beta_{a}(1-p)\sum_{r\not\in A}w_{ar}^{2}(e_{ar}(1)-c_{ar})e_{ar}(1)}_{\text{constant}}$
is the product of a constant (by Lemma 1) and a mean-zero random variable, so $X$, its expectation, is zero.
@@
$\mathbb{E}\left[\widehat{\mathcal{W}_{a}(\mathbf{1})}\mid T_{A}\right]$ $=\underbrace{\sum_{r\in A}T_{r}w_{ar}(e_{ar}-c_{ar})/p}_{=0}$
$\quad{}+\sum_{r\in A}w_{ar}c_{ar}+\sum_{r\not\in A}w_{ar}e_{ar}(1),$
where the first term is zero because $T_{r}w_{ar}(e_{ar}-c_{ar})=T_{r}w_{ar}(e_{ar}(1)-c_{ar})$, and $e_{ar}(1)=c_{ar}=1$ for $r\in A$. This makes $\mathbb{E}\left[\widehat{\mathcal{W}_{a}(\mathbf{1})}\mid T_{A}\right]$ a constant, and thus $Y=0$. ∎
@@
$=\sum_{r\in V_{a}}\left(\frac{T_{r}w_{ar}(e_{ar}(1)-1)}{p}+w_{ar}\right)$
$\quad{}+\sum_{r\not\in V_{a}}\left(\frac{T_{r}w_{ar}e_{ar}(1)}{p}\right)$
@@
$=\sum_{r\in V_{a}}w_{ar}+\sum_{r\in R_{a}\setminus V_{a}}\left(\frac{T_{r}w_{ar}e_{ar}(1)}{p}\right)$
$\quad{}+\sum_{r\not\in R_{a}}\left(\frac{T_{r}w_{ar}e_{ar}(1)}{p}\right)$
@@
[…]
$=\frac{1}{n_{a}^{2}}\sum_{i\in\mathcal{A}}\sum_{j\in\mathcal{I}(i)}\text{Cov}(\hat{\mu}_{i},\hat{\mu}_{j})$
($\hat{\mu}_{i}$ is independent of $\hat{\mu}_{j}$ if $R_{i}\cap R_{j}=\emptyset$)
@@
$|a_{i}|\leq|\hat{\mu}_{i}^{u,c}|+|\mu_{i}|=|\hat{\mu}_{i}^{u,c}|+|\mathbb{E}[\hat{\mu}_{i}^{u,c}]|\leq|\hat{\mu}_{i}^{u,c}|+\mathbb{E}[|\hat{\mu}_{i}^{u,c}|]\\ \leq\frac{2MW_{h}|R_{a}|}{p^{2}(1-p)W_{l}}\leq\frac{2MW_{h}d_{a}}{p^{2}(1-p)W_{l}}.$ (12)
@@
$d_{W}\left(\frac{\hat{\mu}^{u,c}-\mu}{\sqrt{\text{Var}(\hat{\mu}^{u,c})}},Z\right)$
$\leq\frac{d_{a}^{2}d_{r}^{2}}{\sigma^{3}n_{a}^{2}}\left[\frac{2MW_{h}d_{a}}{p^{2}(1-p)W_{l}}\right]^{3}+\sqrt{\frac{28}{\pi}}\frac{d_{a}^{3/2}d_{r}^{3/2}}{\sigma^{2}n_{a}^{3/2}}\left[\frac{2MW_{h}d_{a}}{p^{2}(1-p)W_{l}}\right]^{2}$
$=O\left(\frac{d_{a}^{5}d_{r}^{2}}{\sigma^{3}n_{a}^{2}}+\frac{d_{a}^{7/2}d_{r}^{3/2}}{\sigma^{2}n_{a}^{3/2}}\right)=O\left(\frac{d_{a}^{5}d_{r}^{2}}{n_{a}^{1/2}}+\frac{d_{a}^{7/2}d_{r}^{3/2}}{n_{a}^{1/2}}\right)$
$=O\left(\frac{d_{a}^{5}d_{r}^{2}}{n_{a}^{1/2}}\right).$
@@
Lemma 4 (Lemma 3.6 in Ross (2011)).
@@
We suppress the superscript $(u,c)$ in $\tilde{\mu}$. The validity of test follows by definition. We begin by showing that $c_{\alpha}\to 0$ and $\tilde{\mu}-\mu^{*}\overset{p}{\to}0$, where $\mu^{*}=\frac{1}{n_{a}}\sum_{a}\beta_{a}\sum_{r}(w_{ar}c_{ar})$.
@@
$=\frac{1}{n_{a}^{2}}\sum_{i\in\mathcal{A}}\sum_{j\in\mathcal{J}(i)}\text{Cov}(\tilde{\mu}_{i},\tilde{\mu}_{j}\mid y)$
(if $V_{i}\cap V_{j}=\emptyset$, $z_{i}^{u}$ is independent of $z_{j}^{u}$
and thus $\tilde{\mu}_{i}$ is independent of $\tilde{\mu}_{j}$)
@@
$=\mathbb{E}[\tilde{\mu}^{2}\mid y,|\tilde{\mu}|\geq c_{\alpha}]P(|\tilde{\mu}|\geq c_{\alpha}\mid y)$
$\quad+\mathbb{E}[\tilde{\mu}^{2}\mid y,|\tilde{\mu}|<c_{\alpha}]P(|\tilde{\mu}|<c_{\alpha}\mid y)$
@@
Next, we show $\tilde{\mu}-\mu^{*}\overset{p}{\to}0$. From the fact that $\hat{\beta}_{a}$ is unbiased for $\beta_{a}$, we get
@@
[…]
$=\frac{1}{n_{a}^{2}}\sum_{i\in\mathcal{A}}\sum_{j\in\mathcal{I}(i)}\text{Cov}(\tilde{\mu}_{i},\tilde{\mu}_{j})$
(if $R_{i}\cap R_{j}=\emptyset$, $\tilde{\mu}_{i}$ is independent of $\tilde{\mu}_{j}$)
@@
where $\mathcal{I}(i)=\{j\in\mathcal{R}\mid R_{i}\cap R_{j}\neq\emptyset\}$ is defined at the beginning of the proof of Theorem 2. Now we have shown $\mathbb{E}[\tilde{\mu}-\mu^{*}]=0$ and $\text{Var}(\tilde{\mu}-\mu^{*})\to 0$, which means $\tilde{\mu}-\mu^{*}\overset{p}{\to}0$.